OpenStudy (anonymous):

What is the solution set for |2x – 3| = 13?

jimthompson5910 (jim_thompson5910):

|2x – 3| = 13 2x – 3 = 13 or 2x – 3 = -13 2x = 13 + 3 or 2x = -13 + 3 2x = 16 or 2x = -10 x = 16/2 or x = -10/2 x = 8 or x = -5

OpenStudy (anonymous):

|2x – 3| = 13 ===> 2x-3=13==>x=8 or 2x-3=-13 ===> x=-5